OpenStudy (anonymous):

find an equation of the line having the given slope and containing the given point m=2/3, (7,-3)

OpenStudy (anonymous):

y=mx+b -3=2/3(7)+b b=1.69 close calculate yur self and y=(2/3)x+1.69

OpenStudy (anonymous):

m=2/3.(7.-3) make the line: Y=(2/3)x+b through(7.-3) so -3=14/3 +b b=-3-14/3=-23/3 so y=(2/3)x-23/3